OpenStudy (miaaziz):

im stuck at this question. how to evaluate lim┬(x→∞)⁡〖(cos⁡(2/x))^(x^2 ) 〗

OpenStudy (danjs):

\[\large \lim_{x \rightarrow \infty} [\cos(\frac{ 2 }{ x })]^{x^2}\]

OpenStudy (danjs):

maybe try a natural log to move the exponent...e^ln(x) = x \[\large \lim_{x \rightarrow \infty} [e ^{\ln(\cos(2/x)^{x^2}}]\] \[\large \lim_{x \rightarrow \infty} [e ^{x^2*\ln(\cos(2/x)}]\]

OpenStudy (jango_in_dtown):

here is the sorry. oops it is e^(-2).

OpenStudy (jango_in_dtown):

@Miaaziz

OpenStudy (miaaziz):

thank you @DanJS . thank you @jango_IN_DTOWN